How do you make one of these equations equal 48? The equations and numbers are x+(x+2)+(x+4)=48; 16,17,18
x+(x+1)+(x+2)=48; 15,16,17

Answer by ewatrrr(24785)

 
Hi,
x+(x+2)+(x+4)=48
solving:
3x + 6 = 48
3x = 42
x = 14
x = 14, (x+2) = 16 and (x+4) = 18
14 + 16 + 18 = 48
x+(x+1)+(x+2)=48
solving:
Yes! x = 15
x = 15, (x+1) = 16 and (x+2) = 17
15 + 16 + 17 = 48